OpenStudy (anonymous):

Two equal rectangular sand boxes having one side in common and a square floor each, are made up of 96m^2 of material. What are the dimensions of each box if the volume is to be maximized?

OpenStudy (anonymous):

my solution: Area=8x^2+8xy y=(96-8x^2)/8x V=(4x^2)(y) =(4x^2)((96-8x^2)/8x) V(x)=(96x-8x^3)/2 V'(x)=(96-24x^2)/2 x=2---> it's wrong y=(96-8(2)^2)/8(2) =4 the correct answer is 4 x 4 x 16/7

OpenStudy (anonymous):

@hero @thomaster @hartnn

hartnn (hartnn):

Area=8x^2+8xy how ? a box has 6 faces, so 2 box will have 12 faces, not 16

hartnn (hartnn):

A=4x^2+8xy

OpenStudy (anonymous):

hmm.. you're right.. maybe i got confused with my drawing :/ ok,, let me solve it again..thanks!

hartnn (hartnn):

and how are you calculating the volume ?

OpenStudy (anonymous):

sorry I respond TOO late, i'm doing this question now.. uhm.. i just wanted to ask some questions, like for the area, don't you think it should be A= A=4x^2+7xy because the problem says "Two equal rectangular sand boxes having ONE SIDE IN COMMON" and for the volume, I really messed up with my first solution.. now I realized that the volume should be (2x^2)y, am i right??

OpenStudy (anonymous):

I already got the right answer (^_^)!! since the problem says "Two equal rectangular sand boxes having ONE SIDE IN COMMON and A SQUARE FLOOR each", i finally understand that the formula for the area must be A=2x^2+7xy.. I used the formula for the area as my constraint.. then i substitute it to the volume formula which is V=(2x^2)y then I got the right answer which is 4 x 4 x 16/7 (^_^)